# A Hybrid Algorithm for Systems of Non-interacting Particles

### v2

@@ Abstract
Our focus is on simulating the dynamics of non-interacting particles, which, under certain assumptions, can be formally described by the Dean-Kawasaki equation. The Dean-Kawasaki equation can be solved numerically using standard finite volume methods. However, the numerical approximation implicitly requires a sufficiently large number of particles to ensure the positivity of the solution and accurate approximation of the stochastic flux. To address this challenge, we extend hybrid algorithms for particle systems to scenarios where the density is low. The aim is to create a hybrid algorithm that switches from a finite volume discretization to a particle-based method when the particle density falls below a certain threshold. We develop criteria for determining this threshold by comparing higher-order statistics obtained from the finite volume method with particle simulations. We then demonstrate the use of the resulting criteria for dynamic adaptation in both two- and three-dimensional spatial settings.
@@ 1 Introduction
The dynamics of a system of non-interacting random-walker particles can, in a suitable mathematical setting, be formally described by the Dean-Kawasaki equation. The microscopic model that we are considering is a random walk of independent indistinguishable Brownian particles on a -dimensional torus or with Dirichlet or homogeneous Neumann boundary conditions on a rectilinear domain. Let be the empirical density that is defined by
@@
where is a Dirac distribution. Utilizing the Itô’s formula one can formally derive the equation for , see [10, 19]. The formal form of the equation is the so-called Dean-Kawasaki equation. In the case of the independent particles, this is a stochastic partial differential equation of the form
@@
where is a vector-valued space-time white noise. Note that in the case of interacting particles, the interaction term also appears in the equation. It is well-known [20] that due to divergence operator, square root and irregularity of the noise, this equation is highly singular. Analytically, the only martingale solutions take the form of an atomic measure. Furthermore, such solutions only exist for . Unfortunately, this formal mathematical description is not particularly suitable for direct computation. Nevertheless, standard numerical methods applied to the Dean-Kawasaki equation with truncated noise can be used to approximate integral averages of the solution.
@@
In spite of the theoretical difficulties, interest in the analysis and simulation of Dean-Kawasaki-type equations has significantly increased over the past decade, with applications ranging from social dynamics [13, 18] to fluid dynamics [15]. Equations of the form of Eq. (2) also arise in the context of Dynamic Density Functional Theory (DDFT). A detailed discussion on the relationship between deterministic and stochastic DDFT can be found in [3, 16]. In particular, the advantage of using the stochastic partial differential equation approach over discrete particle methods is clear when examining systems with a large number of particles.
@@
Numerical schemes for Dean–Kawasaki-type equations were considered in [7, 8, 21]. In [7] the authors show that a system of independent Brownian motions can be approximated in an appropriate metric to arbitrary order by a finite difference or finite element discretization of the Dean-Kawasaki SPDE, under the assumption that the initial particle distribution must be bounded away from zero. They do not prove positivity of the approximation and allow for the solution becoming negative. Their error bound includes a term accounting for the solution becoming negative, which happens rarely under the assumption . They also note that in the opposite regime, where the direct simulation of particles would be less expensive than the approximation of the Dean–Kawasaki equation.
@@
A hybrid algorithm for particles systems was first proposed by Alexander et al. [1, 2]. Their work demonstrates the conservative coupling between a stochastic partial differential equation (SPDE) and a particle algorithm in one dimension. They show that the resulting hybrid method can capture the mean and variance accurately for both open and close systems. However, their focus is on systems in which the particle density is sufficiently large that the SPDE description is accurate. Here, we first generalize their approach to multiple spatial dimensions with dynamic adaptation. We then categorize the behavior of the dynamics of higher-order statistics as the particle density becomes small. This categorization provides the criterion for determining where the particle description is required.
The paper is structured as follows. In Section 2.1 we briefly recall the well-posedness results for the Dean-Kawasaki equation and its approximation and introduce the linearized Gaussian approximation. Section 2.2 introduces the finite volume approximation of the Dean-Kawasaki equation and the linearized Gaussian approximation. In Section 3 we introduce the hybrid algorithm in one dimension and describe the adaptive hybrid method in higher dimensions. Section 4 contains computational examples.
@@ 2.1 Well-posedness of the equation and its approximation
As mentioned in the introduction, due to singularity of the Dean-Kawasaki equation, well-posedness results are difficult to obtain. In general, some regularization of the noise is needed, since in [20] the authors showed that the only martingale solution of the Dean-Kawasaki equation are of the form of an empirical measure for independent Brownian motions given by (1). Besides the martingale nature of the solution, they also showed that the equation has no solutions unless is a positive integer and therefore it is potentially not well-suited for numerical analysis and computations. Other solution concepts, such as the stochastic kinetic solution, are also explored. However, since we are interested in computational aspects, this approach will not be addressed here; for more detailed analysis, please refer to [1].
In order to address the sensitivity of the solution concept of the original Dean-Kawasaki equation, in [12] the authors suggest a regularized non-linear SPDE approximation of (2). The modification is constructed to handle situations where the solution approaches zero by replacing the non-Lipschitz square root with a Lipschitz function on the interval . Additionally, the noise is approximated by its ultraviolet cutoff at frequencies of order .
It is proved that, under a suitable coercivity assumption—affecting the choice of parameters and for a given —the modified equation is well-posed in the standard weak PDE sense, making it well-suited for numerical computation. Moreover, it can be shown that the solution remains non-negative, conserves mass, and, with an appropriate selection of parameters, error bounds relative to the original particle system can be established.
To assess how the modified equation differs from the Dean-Kawasaki equation with truncated noise, we note that the modification of the square root becomes relevant only when the solution is small, corresponding to a very low-density regime. In the hybrid method, low-density regimes will be handled using particle dynamics. Therefore, the SPDE that we aim to discretize using the finite volume method (FVM) is the Dean-Kawasaki equation with regularized noise:
@@
where . It is well-known that Gaussian approximation correctly describes to first-order the central limit fluctuations. However, the Gaussian approximation fails to predict the large deviation action functional for the particle model, which is accurately predicted by the Dean-Kawaski equation. Thus, the Gaussian approximation cannot accurately determine the likelihood of observing large deviations from typical hydrodynamic limit behaviour [11].
In [7] the authors compared simulations of the Dean-Kawasaki equation and the linearized Gaussian equation. They observed that two models exhibit similar behaviour for the second moments but diverge for the higher moments, which they expected based on the analogy with [7, Lemma 10]. As noted above, this work assumed a sufficient number of particles per cell, while here we want to address exactly the case when the number of particles per cell can be arbitrary small locally.
@@ 2.2 Numerics
In this section we present a simple finite volume discretization of the Dean-Kawasaki equation in three dimensions. The choice of a finite volume approach is motivated by the construction of the hybrid algorithm and is well-suited to the conservation form of the equation. We also consider a finite volume discretization of the linearized Gaussian approximation given by Eqs. (4-5). For simplicity, we will consider discretization in time using an Euler-Maruyama scheme; investigation of other temporal integrators will be addressed in future work. Without loss of generality, we will assume the volume of the entire domain is one.
@@ 2.2.1 Finite volume discretization
We discretize on a domain of size on a uniform finite volume mesh of size with being cells with mesh spacing , , . We then let
@@
where , etc.
@@
where is Gaussian (normal) distributed random variable with zero mean and unit variance and is the averaging operator defined by
with other stochastic fluxes defined analogously. We note that we can also define deterministic fluxes
@@
Using ’s we then define Gaussian stochastic fluxes by
We can then define the discretizaiton of the inearized Gaussian approximation as
@@ 3 Hybrid discretization
The Dean-Kawasaki (2) equation describes the evolution of a large number of independent Brownian particles. From a computational perspective, when the number of particles in each computational cell is sufficiently large, the stochastic continuum discretization Eq. (7) provides an accurate description of the dynamics. However, when the number of particles per cell becomes sufficiently small, the approximation is no longer accurate. Implicit in Eq. (7) and (11) is an assumption that the fluctuations in the fluxes can be modeled as white noise.
The main goal here is to develop a hybrid algorithm that overcomes the restriction of requiring enough particles per cell. The basic idea of the hybrid algorithm is to apply a particle-based method in the regions of the domain where the number density approximation becomes very small or in the regions where a higher fidelity particle representation is required. The construction of the hybrid algorithm will be based on the adaptive mesh and algorithm refinement (AMAR) approach introduced in [17] and extended to stochastic systems in [4, 14]. Adopting this type of approach, we can treat regions where the particle density is very small (or zero) accurately and ensure positivity of the solution. The algorithm here generalizes the one-dimensional algorithm in [1] to multiple spatial dimensions and incorporates dynamic refinement.
@@
Constructing an AMAR hybrid involves a number of elements. We need to define how to map between different representations of the solution. We need to define a synchronization procedure that corrects results of advancing the SPDE and particle regions sequentially to obtain a composite solution. Finally, we need to determine where the finer-scale model is required; i.e., when the SPDE model does not provide a sufficiently accurate representation of the system. One potential criterion here would be simply maintaining positivity of the solution. However, this assumes that the SPDE representation is adequate up until the point when the solution becomes negative. We will explore this issue in more detail and suggest alternative criteria in the next section.
The mapping from the particle representation to the SPDE representation is straightforward. We view the particles as point particles so we can define the SPDE representation as the number of particles in the cell divided by the cell volume. The mapping from the SPDE representation to a particle representation is not unique. Here we will view the problem of mapping from SPDE to particles as a conditional sampling problem, namely, we want to generate a random set of particles that are, at least approximately, consistent with the SPDE representation. Given the SPDE description we can compute the number of particle in each cell. Those particle are then assigned random locations within the cell where we assume that the particles are uniformly distributed in the cell. We note that, in general, the number of particles in a cell may not be an integer. If this occurs we use a probabilistic approach to define the particles. Specifically, if a cell contains particles where and then we first insert particles as described above. We then choose a random number from the uniform distribution on . If then we insert an additional particle; otherwise we do not. This procedure preserves the expected value of the number of particles in the system but is not strictly conservative.
@@ 3.1 Hybrid algorithm in one dimension
We are now ready to discuss the time-step algorithm for the hybrid. For simplicity of the exposition, we will first describe the algorithm in one dimension and assume that the particle region is a contiguous collection of cells, to that does not change with time. We will then describe the steps needed to extend the overall approach to multiple dimension with dynamic particle patches that consists of different (not necessarily) connected cells. We will also assume that the SPDE and the particle components of the algorithm take the same time step; however, this restriction can also be relaxed. We assume that the system has been initialized by defining the location of each particle in the particle region and by specifying the number density in the SPDE region outside the particle region. We note that using the mappings defined above, we can initialize the system from either a pure particle description or from an SPDE description. As noted above, in AMAR the coarse, SPDE representation is maintained throughout the entire domain. Thus, after initializing the particle representation we map the particle solution onto the SPDE solution in the region covered by the particles.
@@
To evolve the system, we first advance the SPDE for a timestep over the entire domain using Eq. (7) and denote the result by . To advance the particle region, we need to supply suitable boundary conditions. The boundary conditions are necessary to capture the effect of particles entering the particle region during a time step. These boundary conditions are supplied by the SPDE solution. In the cells adjacent to the particle region, and we compute the number of particles in each cell from the SPDE solution. We then populate these boundary cells with the specified number of particles with random locations within the cell. We then advance the particle region by allowing particles in the region to perform a random walk step. (We note here that we limit the motion of a single particle to be at most so that a particle can move at most one cell in a time step. We select the time step so that violating this restriction would be extremely rare.) We then define the new particle representation to be the collection of particle in cells to . (Particles outside the particle region at the end of the step are discarded.)
@@
We now need to synchronize the two representations. The idea here is that the particle representation is higher fidelity than the SPDE representation so the composite solution should reflect the particle information. There are two different components to the synchronization. First, for SPDE cells that are within the particle region , we replace the SPDE solution with the value obtained from counting the number of particles in the cell, which we denote by .
@@
where is the net number of particle that crossed edge from left to right during the time step. After these adjustments, the SPDE values at and have been effectively updated using the particle flux instead of the SPDE flux. For cells away from the particle region we simply set to complete the synchronization. The combination of supplying boundary conditions for the particles from the SPDE solution and synchronizing the fluxes using Eq.(12) effectively creates a composite solution of the system. The process is analogous to solving a second-order parabolic PDE in different subdomains and combining them by matching Dirichlet and Neumann boundary conditions at the interface between the domains.
We note that here we have used the same time step for the SPDE and the particles, where the time step was selected so that a particle moving more than during a time step is a rare event. A smaller time step can be used in the particle region where the particles in the boundary cells for intermediate steps are computed from interpolation of the SPDE solution in time and the number of particles crossing the boundary is summed over the substeps.

### v3

@@ Abstract
Our focus is on simulating the dynamics of non-interacting particles including the effects of an external potential, which, under certain assumptions, can be formally described by the Dean-Kawasaki equation. The Dean-Kawasaki equation can be solved numerically using standard finite volume methods. However, the numerical approximation implicitly requires a sufficiently large number of particles to ensure the positivity of the solution and accurate approximation of the stochastic flux. To address this challenge, we extend hybrid algorithms for particle systems to scenarios where the density is low. The aim is to create a hybrid algorithm that switches from a finite volume discretization to a particle-based method when the particle density falls below a certain threshold. We develop criteria for determining this threshold by comparing higher-order statistics obtained from the finite volume method with particle simulations. We then demonstrate the use of the resulting criteria for dynamic adaptation in both two- and three-dimensional spatial settings in the absence of an external potential. Finally we consider the dynamics when an external potential is included.
@@ 1 Introduction
The dynamics of a system of non-interacting random-walker particles under the influence of an external potential can, in a suitable mathematical setting, be formally described by the Dean-Kawasaki equation.
@@
Specifically the microscopic model that we are considering is a random walk of independent indistinguishable Brownian particles under the influence of an external potential on a -dimensional torus or with Dirichlet or homogeneous Neumann boundary conditions on a rectilinear domain. We assume that the particle positions, , evolve according to a Langevin equation
We let be the empirical density defined by
@@
where is a Dirac distribution. Utilizing the Itô’s formula one can formally derive the equation for , see [13, 29, 8]. The formal form of the equation is the so-called Dean-Kawasaki equation. In the case of independent particles including an external potential, this is a stochastic partial differential equation (SPDE) of the form
@@
where is a vector-valued space-time white noise. Note that in the case of interacting particles, the interaction term also appears in the equation. It is well-known [30] that due to divergence operator, square root and irregularity of the noise, this equation is highly singular, meaning that the standard SPDE theory for singular SPDEs [24, 23] is not directly applicable. Analytically, the only martingale solutions of Eq. (3) take the form of an atomic measure. Furthermore, such solutions only exist for . This implies that, from a rigorous mathematical perspective, solutions to Eq. (3) can only be determined by directly modeling the individual particles. A common approach to circumvent these theoretical issues is to introduce an high-frequency cutoff for the noise, which then gives a regularized version of Eq. (3)
where represents the regularized noise. We note that in this form, is no longer required to be an integer.
@@
The advantage of using the SPDE approach over discrete particle methods is clear when examining systems with a large number of particles. Interest in the analysis and simulation of these types of regularized Dean-Kawasaki equations has significantly increased over the past decade, with applications ranging from social dynamics [16, 26] to fluid dynamics [18]. Equations of the form of Eq. (3) also arise in the context of Dynamic Density Functional Theory (DDFT). A detailed discussion on the relationship between deterministic and stochastic DDFT can be found in [3, 19].
@@
Numerical schemes for Dean–Kawasaki-type equations were considered in [10, 11, 32]. In [10] the authors show that a system of independent Brownian motions can be approximated in an appropriate metric to arbitrary order by a finite difference or finite element discretization of the Dean-Kawasaki SPDE, when the system contains a sufficiently large number of particles and the initial distribution is bounded away from zero. They do not prove positivity of the approximation and allow for the solution becoming negative. Their error bound includes a term accounting for the solution becoming negative, which happens rarely under the assumption . They also note that in the low particle regime where , which corresponds to less than or equal to one particle per mesh cell, the direct simulation of particles would be less expensive than approximation of the Dean–Kawasaki equation.
@@
A hybrid algorithm for particles systems was first proposed by Alexander et al. [1, 2]. Their work demonstrates the conservative coupling between an SPDE and a particle algorithm in one dimension. They show that the resulting hybrid method can capture the mean and variance accurately for both open and close systems. However, their focus is on systems in which the particle density is sufficiently large that the SPDE description is accurate. Here, we first generalize their approach to multiple spatial dimensions with dynamic adaptation. We then categorize the behavior of the dynamics of higher-order statistics as the particle density becomes small. This categorization provides the criterion for determining where the particle description is required.
The paper is structured as follows. In Section 2.1 we briefly recall the well-posedness results for the Dean-Kawasaki equation and its approximation and introduce the linearized Gaussian approximation. Section 2.2 introduces the finite volume approximation of the Dean-Kawasaki equation and the linearized Gaussian approximation. In Section 3 we introduce the hybrid algorithm in one dimension and describe the adaptive hybrid method in higher dimensions. Section 4 contains computational examples that serve to validate the algorithm in the absence of an external potential and illustrate its capabilities when an external potential is included.
@@ 2.1 Well-posedness of the equation and its approximation
As mentioned in the introduction, due to singularity of the Dean-Kawasaki equation, well-posedness results are difficult to obtain. In general, some regularization of the noise is needed, since in [30] the authors showed that the only martingale solution of the Dean-Kawasaki equation is of the form of an empirical measure for independent Brownian motions given by (2). Besides the martingale nature of the solution, they also showed that the equation has no solutions unless is a positive integer and therefore it is potentially not well-suited for numerical analysis or computation. Other solution concepts, such as the stochastic kinetic solution, are also explored. However, since we are interested in computational aspects, this approach will not be addressed here; for more detailed analysis, please refer to [20]. .
In order to address the sensitivity of the solution concept of the original Dean-Kawasaki equation, in [15] the authors suggest a regularized non-linear SPDE approximation of (3). The modification is constructed to handle situations where the solution approaches zero by replacing the non-Lipschitz square root with a Lipschitz function that regularizes it around zero. Additionally, the noise is approximated by introducing an ultraviolet cutoff for frequencies above order .
The authors prove that, under a suitable coercivity assumption, which relates the regularization of the square root and the noise truncation for a given , the modified equation is well-posed in the standard weak PDE sense, making it well-suited for numerical computation. Moreover, it can be shown that the solution remains non-negative, conserves mass, and, with an appropriate selection of parameters, error bounds relative to the original particle system can be established.
To assess how the modified equation differs from the Dean-Kawasaki equation with truncated noise, we note that the modification of the square root becomes relevant only when the solution is small, corresponding to a very low-density regime. In the hybrid method, low-density regimes will be handled using particle dynamics. Thus for the hybrid method, we aim to discretize the unmodified SPDE Eq. (4) using a finite volume methods (FVM).
@@
where . It is well-known that the Gaussian approximation correctly describes to first-order the central limit fluctuations. However, the Gaussian approximation fails to predict the large deviation action functional for the particle model, which is accurately predicted by the Dean-Kawaski equation. Thus, the Gaussian approximation cannot accurately determine the likelihood of observing large deviations from typical hydrodynamic limit behaviour [14].
In [10] the authors compared simulations of the Dean-Kawasaki equation and the linearized Gaussian equation. They observed that two models exhibit similar behaviour for the second moments but diverge for the higher moments, which they expected based on the analogy with [10, Lemma 10]. As noted above, this work assumed a sufficient number of particles per cell, while here we want to address exactly the case when the number of particles per cell can be arbitrary small locally.
@@ 2.2 Numerics
In this section we present a simple finite volume discretization of the Dean-Kawasaki equation in three dimensions. For simplicity of the notation, we will present the basic methodology without including an external potential. The modifications needed to include an external potential can be found in Appendix A. The choice of a finite volume approach is motivated by the construction of the hybrid algorithm and is well-suited to the conservation form of the equation. We also consider a finite volume discretization of the linearized Gaussian approximation given by Eqs. (5-6). For simplicity, we will consider discretization in time using an Euler-Maruyama scheme; investigation of other temporal integrators will be addressed in future work. Without loss of generality, we will assume the volume of the entire domain is one.
@@ 2.2.1 Finite volume discretization
We discretize on a domain of size on a uniform finite volume mesh of size with mesh spacing , , . We then let
@@
where , etc. The structure of the mesh in 2D is illustrated in Figure 1.
@@
where is a Gaussian (normal) distributed random variable with zero mean and unit variance associated with edge and is the averaging operator defined by
with other stochastic fluxes defined analogously. This stochastic flux approximates the regularized noise term in Eq. (4) where the regularization is given by the mesh spacing. We note that we can also define deterministic fluxes
@@
Using the ’s, we define Gaussian stochastic fluxes by
We can then define the discretization of the inearized Gaussian approximation as
@@ 3 Hybrid discretization
The Dean-Kawasaki (3) equation describes the evolution of a large number of independent Brownian particles with an external potential. From a computational perspective, when the number of particles in each computational cell is sufficiently large, the stochastic continuum discretization Eq. (8) provides an accurate description of the dynamics. However, when the number of particles per cell becomes sufficiently small, the approximation is no longer accurate. Implicit in Eq. (8) and (12) is an assumption that the fluctuations in the fluxes can be modeled as white noise.
The main goal here is to develop a hybrid algorithm that overcomes the restriction of requiring a sufficiently large number of particles per cell. The basic idea of the hybrid algorithm is to apply a particle-based method in the regions of the domain where the number density becomes very small or in the regions where a higher fidelity particle representation is required. The construction of the hybrid algorithm will be based on the adaptive mesh and algorithm refinement (AMAR) approach introduced in [22] and extended to stochastic systems in [4, 17]. Adopting this type of approach, we can treat regions where the particle density is very small (or zero) accurately and ensure positivity of the solution. The algorithm here generalizes the one-dimensional algorithm in [1] to multiple spatial dimensions, includes an external potential, and incorporates dynamic refinement.
@@
Constructing an AMAR hybrid involves a number of elements. We need to define how to map between different representations of the solution. We need to define a synchronization procedure that corrects results of advancing the SPDE and particle regions sequentially to obtain a composite solution. Finally, we need to determine where the finer-scale model is required; i.e., when the SPDE model does not provide a sufficiently accurate representation of the system. One potential criterion here would be simply maintaining positivity of the solution. However, this assumes that the SPDE representation is adequate up until the point when the solution becomes negative. We will explore this issue in more detail and suggest alternative criteria in the next section. The mapping from the particle representation to the SPDE representation is straightforward. We view the particles as point particles so we can define the SPDE representation as the number of particles in the cell divided by the cell volume. The mapping from the SPDE representation to a particle representation is not unique. Here we will view the problem of mapping from SPDE to particles as a conditional sampling problem, namely, we want to generate a random set of particles that are, at least approximately, consistent with the SPDE representation. Given the SPDE description we can compute the number of particle in each cell. Those particle are then assigned random locations within the cell where we assume that the particles are uniformly distributed in the cell for the case when there is no external potential. When an external potential is included the sampling needs to be modified, as discussed in Appendix A. We note that, in general, the number of particles in a cell may not be an integer. If this occurs we use a probabilistic approach to define the particles. Specifically, if a cell contains particles where and then we first insert particles as described above. We then choose a random number from the uniform distribution on . If then we insert an additional particle; otherwise we do not. This procedure preserves the expected value of the number of particles in the system but is not strictly conservative.
@@ 3.1 Hybrid algorithm in one dimension
We are now ready to discuss the time-step algorithm for the hybrid. For simplicity of the exposition, we will first describe the algorithm in one dimension and assume that the particle region is a contiguous collection of cells from to that does not change with time. We will then describe the steps needed to extend the overall approach to multiple dimension with dynamic particle patches that consist of different (not necessarily) connected cells. We will also assume that the SPDE and the particle components of the algorithm take the same time step; however, this restriction can also be relaxed. We assume that the system has been initialized by defining the location of each particle in the particle region and by specifying the number density in the SPDE region outside the particle region. We note that using the mappings defined above, we can initialize the system from either a pure particle description or from an SPDE description. As noted above, in AMAR the coarse, SPDE representation is maintained throughout the entire domain. Thus, after initializing the particle representation we map the particle solution onto the SPDE solution in the region covered by the particles.
@@
To evolve the system, we first advance the SPDE for a timestep over the entire domain using Eq. (8) and denote the result by . To advance the particle region, we need to supply suitable boundary conditions. The boundary conditions are necessary to capture the effect of particles entering the particle region during a time step. These boundary conditions are supplied by the SPDE solution. In the cells adjacent to the particle region, and we compute the number of particles in each cell from the SPDE solution. We then populate these boundary cells with the specified number of particles with random locations within the cell. We then advance the particle region by allowing particles in the region to perform a random walk step. Here, we limit the motion of a single particle to be at most so that we do not need to account for particles moving more than one cell. For the time steps used here, is sufficiently small that a particle moving more than in a time step would be a rare event. Specifically for the parameters used here (see Sec. 4) the probability of a step larger than is less than . Furthermore, empirically we find that when a step length is reduced the reduction is typically less than . We have found no evidence that these small adjustments in rare cases impact the statistics of the simulations. (We note that one can enlarge the domain of influence of a particle to allow larger particle steps by enlarging the regions where the SPDE provides boundary data for the particle region.) We then define the new particle representation to be the collection of particle in cells to . (Particles outside the particle region at the end of the step are discarded.)
@@
We now need to synchronize the two representations. The idea here is that the particle representation is higher fidelity than the SPDE representation so the composite solution should reflect the particle information. There are two different components to the synchronization. First, for SPDE cells that are within the particle region , we replace the SPDE solution with the value obtained from counting the number of particles in the cell and dividing by the cell volume, which we denote by .
@@
where is the net number of particle that crossed edge from left to right during the time step. After these adjustments, the SPDE values at and have been effectively updated using the particle flux instead of the SPDE flux. For cells away from the particle region we simply set to complete the synchronization. The combination of supplying boundary conditions for the particles from the SPDE solution and synchronizing the fluxes using Eq.(13) effectively creates a composite solution of the system. We note that after the completion of this step, the hybrid integration step is conservative. The process is analogous to solving a second-order parabolic PDE in different subdomains and combining them by matching Dirichlet and Neumann boundary conditions at the interface between the domains.
We note that here we have used the same time step for the SPDE and the particles. A smaller time step can be used in the particle region where the particles in the boundary cells for intermediate steps are populated using the SPDE solution interpolated temporally to the intermediate times and the number of particles crossing the boundary is summed over the substeps.